The problem is y^2-6y+16x+25=0. I have to find the vertex, the focus, the directrix and the axis.
Answer by stanbon(75887)

The problem is y^2-6y+16x+25=0.
I have to find the vertex, the focus, the directrix and the axis.
-----------------------------
Rearrange:
16x = -y^2+6y-25
16x = -(y^2-6y+25)
16x = -(y^2-6y+9)-25+9
--------
x = (-1/16)(y-3)^2 +(9/16)
That is a parabola opening to the left:
k = 3 ; h = 9/16
4p = -1/16
p = -1/64
-----------------
Vertex: (9/16 , 3)
Axis: y = 3
Focus: (9/16-1/64,3) = (35/64 , 3)
Directrix: x = 9/16+1/64 = 37/64
=========================================
Cheers,
Stan H.
=================